Main INDEX
Monthly INDEX
PREV
NEXT
User name lkaufman
Log entry time 23:59:44 on July12,2004
Entry number 128407
keyword=Swing Shift Summary
Swing Shift Summary
Begin shift with no beam to a little bit of beam while accelerator works on the
compton tune.
Put LH2 target back in around 18:50. We also turned the raster on to 12x12
(3x3.5).
18:58 Starting taking data again with 37 uA of beam, raster on with LH2 target.
19:39 Beam down for a while for injector access.
23:25 Update from MCC estimates 1-2 more hours necessary for work on Hall C
laser.
Ended shift with no beam.
Run #
------
2827 - pedestal/compton tuning, empty target
2828 - compton tuning, empty target
2829 - compton tuning, empty target
2830 - compton tuning, empty target
2831 - production, only 45 min with several beam trips
2832 - pedestal run
2833 - pedestal run
2834 - pedestal run
2835 - pedestal run
2836 - pedestal run